In the unit circle pictured below, what is the measurement of angle e? a) 330° Ob) -240 Oc) 210 Od) -210 Oe) -150°
Answer #1

Son Here - B Here consider OABC where OA = 13 & AL - So tana = AB 42 - On tana = I Bh a= tan (4) - 300 Ota - 360° » O= 360°-a